# Jone Lopez de Gamiz Zearra Stevenson Center 1323

### v2

@@ 5. Artin groups of FC-type and parabolic subgroups
Recall that a subgroup $H$ of a group $G$ is weakly malnormal if there is some element $g\in G$ such that $H\cap H^{g}=1$ and an Artin group is irreducible if it does not split as a direct product of standard parabolic subgroups. As a consequence, Charney, Martin and Morris-Wright deduce in [8] that these two properties (that is, being irreducible and having the parabolic intersection property) also imply that the group is acylindrically hyperbolic. In particular, they obtain that irreducible even Artin groups of FC-type are acylindrically hyperbolic. Recall [21, Definition 1.1] that a group $G$ is called acylindrically hyperbolic is there is a possibly infinite generating set $X$ such that the associated Cayley graph $\mathcal{C}$ is hyperbolic with the boundary consisting of more than 2 points and, moreover, the action of $G$ on $\mathcal{C}$ is acylindrical, meaning that for every $\varepsilon>0$ there are $R,N>0$ such that whenever $x,y\in\mathcal{C}$ are such that $d(x,y)\geq R$, then there are at most $N$ elements $g\in G$ with $d(x,gx)\leq\varepsilon$ and $d(y,gy)\leq\varepsilon$.
@@
We first need a couple of technical result about normalizers and centralizers. Given an Artin group $G=A_{\Gamma}$ and a standard parabolic subgroup $A_{S}$, $S\subseteq\Gamma$, Godelle determined in [17, Corollary 0.4] a generating family for the quasi-centralizer, $QC_{G}(A_{S})$, which is the subgroup
@@
As Antolín and Foniqi have observed in [2], if $A_{\Gamma}$ is even, the quasi-centralizer coincides with the centralizer $C_{G}(A_{S})$ and, as a consequence, they prove in [2, Lemma 2.12, Lemma 2.13] the equalities
$N_{G}(A_{S})=A_{S}C_{G}(A_{S})=A_{S}\bigcap_{a\in S}C_{G}(a).$
@@
As we noted above, as $\Gamma$ is even we can use [17, Corollary 0.4] to obtain a generating system for the centralizer of any standard parabolic subgroup $A_{S}$. More explicitely, Godelle proves that there is a generating system consisting on elementary ribbons, which are elements $g$ of a particular type such that $g^{-1}Xg=Y$ where $X,Y\subseteq\Gamma$. The fact that $\Gamma$ is even implies that if there is such an element $g$, then we must have that $X=Y$, since we have the retraction $A_{\Gamma}\to A_{X}$. The elementary ribbons are described in terms of the group elements $\Delta_{X}\in G$ where $X\subseteq\Gamma$ is a spherical subset. We do not need a general description of $\Delta_{X}$, but just the fact that if $X=\{a\}$, then $\Delta_{X}=a$ and if $X=\{a,w\}$ is an edge with label $2k$, then $\Delta_{X}=(wa)^{k}$. By [17, Corollary 0.4], in even Artin groups the elementary ribbons that generate $C_{G}(A_{S})$ are of one of the following two types:
@@
Applying this to the case when $X=\{a\}$, we see that $C_{G}(a)$ is generated by $a$, all those $v\in V$ such that $[v,a]=1$ and an element of the form $(wa)^{k}a^{-1}$ for each $w\in V$ linked with $a$ with label $2k$ for $k>1$. As all these elements lie in $A_{\operatorname{st}_{\Gamma}(a)}$, so does the subgroup $C_{G}(a)$.
@@
$z_{a,w}=a^{1-k}(wa)^{k}a^{-1}=w^{a^{k-1}}w^{a^{k-2}}\ldots w,$
for each $w\in V$ linked with $a$ with label $2k$ for $k>1$.
As $A_{\operatorname{st}_{\Gamma}(a)}$ is parabolic, we have an explicit presentation for $A_{\operatorname{st}_{\Gamma}(a)}$. The generators are the vertices of $\operatorname{st}_{\Gamma}(a)$ and there are two types of relators that involve $a$: some of the form $[v,a]$ and some of the form $(wa)^{k}=(aw)^{k}$ for $k>1$. We can modify this presentation adding generators of the form $w_{i}=w^{a^{i}}$ for $21\leq i\leq k-$ 2 and the elements $z_{a,w}$ as above. Writing $w_{0}$ instead of $w$, we get new relators
$w_{i}^{a}=w_{i+1}$
@@
and furthermore the relator $(wa)^{k}=(aw)^{k}$ yields
@@
The FC-condition implies that if two vertices $w$ and $u$ are linked to $a$ with labels greater than 2, then they cannot be adjacent. If such an $w$ is linked to some $v\in V$ with $[v,a]=1$, then the label must be 2, so we also get that $[w_{i},v]=1$ and $[z_{a,w},v]=1$. With this new presentation we see that $A_{\operatorname{st}_{\Gamma}(a)}$ is isomorphic to a semidirect product $\langle a\rangle M$, where $M$ is an Artin group with vertices the elements $v\in V$ with $[v,a]=1$, and, for each $w$ linked to $a$ with label $2k$, $k>1$, vertices $w_{i}$ and $z_{a,w}$ as before. Let $L$ be the subgroup generated by the elements $v\in V$ with $[v,a]=1$, and, for each $w$ linked to $a$ with label $2k$, $k>1$, the vertex $z_{a,w}$. Then $L$ is an Artin group, and together with $a$, it generates $C_{G}(a)$. Note that $L$ is based precisely on the subgraph $\operatorname{lk}_{\Gamma}(a)$ which is isomorphic to a proper subgraph of $\Gamma$ (but with different vertices).
@@
We have that $C_{G}(a)\leq A_{\operatorname{st}_{\Gamma}(a)}$, so
@@
Consider first the case when $S=\{a\}$ has only one element. Then, as $N_{G}(A_{S})=C_{G}(a)$, using Lemma 5.1 we have (iii).
@@
and using the previous case we get again (iii).
@@
For this, it is enough to show that $C_{G}(A_{S})=C_{G_{1}}(A_{S})$. Consider the subgraph induced by the three vertices $a$, $b$, $v$. Lemma 5.1 implies that $C_{G}(a)\cap C_{G}(b)=C_{G_{1}}(a)\cap C_{G_{1}}(b)$. Therefore,
@@
This implies that we can succesively remove from $\Gamma$ vertices of $\Gamma\setminus S$ without affecting the normalizer until we get a subgraph $S\subseteq\Gamma_{2}\subseteq\Gamma$ such that for any vertex $w$ of $\operatorname{lk}_{\Gamma_{2}}(S)$ and any $a\in V(S)$ so that $w$ and $a$ are linked by an edge of label $2k$ with $k>1$, the vertex $a$ commutes with any other vertex in $S$, that is, it is central in $A_{S}$. Hence, there is some $Z\subseteq S$ (possibly empty) with $A_{Z}$ central in $A_{S}$ and such that all the edges between $S$ and $\operatorname{lk}_{\Gamma_{2}}(S)$ with label greater than 2 have an endpoint in $Z$. As $S$ is not free abelian, $\varnothing\neq S\setminus Z$, and the elements of $A_{S\setminus Z}$ commute with the elements of $A_{\operatorname{lk}_{\Gamma_{2}}(S)}$. Therefore, for $G_{2}=A_{\Gamma_{2}}$ we have that
@@
We follow the steps taken by Minasyan-Osin in [21, Lemma 6.10] where they show the analogous result for graph products of groups.
@@
where we abuse notation and we denote by $S(Y)$ the subgraph of $\Gamma$ induced by $S(Y)$. Since $\Gamma$ is finite, the function, that corresponds to each finite subset $Y\subseteq X$ the integer value $|S(Y)|$, attains its maximum on some finite subset $X^{\prime}\subseteq X$. Hence, to prove the result it is enough to show that $X\subseteq P_{c_{\Gamma}}(X^{\prime})$.
@@
$P_{c_{\Gamma}}(X^{\prime})=P_{c_{\Gamma}}(X).$
The subgroup $H^{\prime}\coloneqq\langle X^{\prime}\rangle\leq H$ is finitely generated and every element of $H^{\prime}$ fixes a vertex of $T$, so $H^{\prime}$ fixes some vertex of $T$. Hence, there is $f\in G$ such that $H^{\prime}\leq fA_{D}f^{-1}$, where $D$ is either $\operatorname{st}_{\Gamma}(w)$ or $\Gamma\setminus\{w\}$. It follows that
@@
    there are Artin groups $A_{T_{1}}$, $A_{T_{2}}$ with $|V(T_{1})|,|V(T_{2})|<|V(\Gamma)|$ such that $H\leq A_{T_{1}}\times A_{T_{2}}$; or
@@
The proof is similar to the one of [21, Corollary 6.19]. Hence, we refer the reader to such paper for a deeper understanding of the main concepts and results.
@@
Let us secondly prove that we can find a hyperbolic element $g\in H$ such that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is normalized by $H$ in $A_{\Gamma}$. Following [21, Definition 6.2], if $P=fA_{S}f^{-1}$ is a parabolic subgroup of $A_{\Gamma}$ for some subgraph $S$ of $\Gamma$ and $f\in A_{\Gamma}$, then the parabolic dimension $\text{pdim}_{\Gamma}(P)$ of $P$ is the number of vertices of $S$. Then [1, Lemma 3.2] ensures that $\text{pdim}_{\Gamma}(P)$ is well-defined, as the result shows that if $g_{1}A_{S}g_{1}^{-1}\subseteq g_{2}A_{T}g_{2}^{-1}$ for some subgraphs $S,T$ of $\Gamma$ and $g_{1},g_{2}\in A_{\Gamma}$, then $S\subseteq T$.
We claim that if $g$ is a hyperbolic element such that $\text{pdim}_{\Gamma}(\text{PStab}_{A_{\Gamma}}(\mathcal{L}))$ is minimal, then $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is normalized by $H$ in $A_{\Gamma}$. This is exactly shown as [21, Proposition 6.13]. The properties used here are proved in [1], and they correspond to [1, Lemma 3.1, Lemma 3.4, Lemma 6.1 (a), Lemma 6.1 (b), Lemma 6.3 (a)].
By Theorem 5.3 there are indeed hyperbolic elements in $H$ for the given action on the Bass-Serre tree $T$. If we consider $h$ an element in $H$ which is hyperbolic and such that $\text{pdim}_{\Gamma}(\text{PStab}_{A_{\Gamma}}(\mathcal{L}))$ is minimal, where $\mathcal{L}$ is the axis of $h$, then we have shown that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is normalized by $H$ and parabolic in $A_{\Gamma}$, so there is a subgraph $S$ of $\Gamma$ and $f\in A_{\Gamma}$ such that $fA_{S}f^{-1}$ is normalized by $H$. But then $H\leq N_{G}(fA_{S}f^{-1})$ and as $H$ is not a subgroup of $fA_{S}f^{-1}$ (otherwise $H$ would fix a point in $T$) by Proposition 5.2, we get (i). Otherwise, we conclude that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is trivial. We have shown that there are two vertices $x$ and $y$ such that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})=\text{PStab}_{A_{\Gamma}}(\{x,y\})$, so [21, Corollary 4.3] implies that $h$ satisfies the WPD condition. Then, from [21, Corollary 2.2], it follows that the action of $H$ on the associated Bass-Serre tree is acylindrically hyperbolic (since we are assuming that $H$ is not virtually cyclic). ∎
@@
If $H$ is a subgroup of $A_{\Delta}$ such that $\Delta$ is a complete graph, then either we can embed $H$ in a decomposition as before or $A_{\Delta}$ is dihedral (recall that $\Delta$ is even of FC-type). If $G=A_{\Delta}$ is dihedral associated to the edge $\{a,b\}$ with label $2k$, the defining relator is $(ab)^{k}=(ba)^{k}$. Setting $c$ to be $ab$, this relator can be transformed into $c^{k}b=bc^{k}$. Then, if we denote by $C$ the subgroup generated by $c^{k}$, we see that $C$ is central in $G$ (in fact, it is the center of $G$). Let $F$ be the normal subgroup of $G$ normally generated by $b$. It is known (see, for example, [2, Lemma 2.5]) that $F$ is free and $C\times F$ has finite index in $G$. Moreover, using the presentation of $G$ in terms of $c$ and $b$ we have that $G/C=\mathbb{Z}\ast\mathbb{Z}_{k}$. Now, if $H\cap C=1$, then $H$ is isomorphic to a subgroup of $G/C$ and by [21, Theorem 2.12] we conclude that $H$ is acylindrically hyperbolic. Thus, we assume that $H\cap C\neq 1$. But as $H$ is not virtually cyclic, $H\cap F\neq 1$ and we are in case (i).
@@
Acylindrically hyperbolic groups cannot contain normal subgroups as in (i) (see [21, Theorem 3.7 (b), (c)] so this result implies that for subgroups of even Artin groups of FC-type this obstruction characterizes acylindrical hyperbolicity.
6. Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
7. Data availability
We do not analyze or generate any datasets, because our work proceeds within a theoretical and mathematical approach. One can obtain the relevant materials from the references below.

### v3

@@ 5. Artin groups of FC-type and parabolic subgroups
Recall that a subgroup $H$ of a group $G$ is weakly malnormal if there is some element $g\in G$ such that $H\cap H^{g}=1$ and an Artin group is irreducible if it does not split as a direct product of standard parabolic subgroups. As a consequence, Charney, Martin and Morris-Wright deduce in [10] that these two properties (that is, being irreducible and having the parabolic intersection property) also imply that the group is acylindrically hyperbolic. In particular, they obtain that irreducible even Artin groups of FC-type are acylindrically hyperbolic. Recall [25, Definition 1.1] that a group $G$ is called acylindrically hyperbolic is there is a possibly infinite generating set $X$ such that the associated Cayley graph $\mathcal{C}$ is hyperbolic with the boundary consisting of more than 2 points and, moreover, the action of $G$ on $\mathcal{C}$ is acylindrical, meaning that for every $\varepsilon>0$ there are $R,N>0$ such that whenever $x,y\in\mathcal{C}$ are such that $d(x,y)\geq R$, then there are at most $N$ elements $g\in G$ with $d(x,gx)\leq\varepsilon$ and $d(y,gy)\leq\varepsilon$.
@@
We first need a couple of technical results about normalizers and centralizers. Given an Artin group $G=A_{\Gamma}$ and a standard parabolic subgroup $A_{S}$, $S\subseteq\Gamma$, Godelle determined in [20, Corollary 0.4] a generating family for the quasi-centralizer, $QC_{G}(A_{S})$, which is the subgroup
@@
As Antolín and Foniqi have observed in [2, Lemma 2.13], if $A_{\Gamma}$ is an even Artin group of FC-type, then the quasi-centralizer coincides with the centralizer $C_{G}(A_{S})$ and, as a consequence, they prove in [2, Theorem 2.12, Lemma 2.13] the equalities
(1) $N_{G}(A_{S})=A_{S}\cdot C_{G}(A_{S})=A_{S}\bigcap_{a\in S}C_{G}(a).$
@@
As we noted above, as $\Gamma$ is even we can use [20, Corollary 0.4] to obtain a generating system for the centralizer of any standard parabolic subgroup $A_{S}$. More explicitely, Godelle proves that there is a generating system consisting on elementary ribbons, which are elements $g$ of a particular type such that $g^{-1}Xg=Y$ where $X,Y\subseteq\Gamma$. The fact that $\Gamma$ is even implies that if there is such an element $g$, then we must have that $X=Y$. To see this, assume that there is some $x\in X$, $x\not\in Y$ such that $g^{-1}xg$ equals $y\in Y$. Applying the retraction $\pi\colon A_{\Gamma}\to A_{Y}$ as at the end of Subsection 2.1, we get that $y=\pi(y)=\pi(g^{-1}xg)=\pi(g^{-1})\pi(g)=1$, which is a contradiction. This shows that $X\subseteq Y$, and by symmetry, we deduce that $X=Y$. The elementary ribbons are described in terms of the group elements $\Delta_{X}\in G$ where $X\subseteq\Gamma$ is a spherical subset. We do not need a general description of $\Delta_{X}$, but just the fact that if $X=\{a\}$, then $\Delta_{X}=a$ and if $X=\{a,w\}$ is an edge with label $2k$, then $\Delta_{X}=(wa)^{k}$. By [20, Corollary 0.4], in even Artin groups the elementary ribbons that generate $C_{G}(A_{S})$ are of one of the following two types:
@@
Applying this to the case when $S=\{a\}$, we see that $C_{G}(a)$ is generated by $a$, all those $v\in V$ such that $[v,a]=1$ and an element of the form $(wa)^{k}a^{-1}$ for each $w\in V$ linked with $a$ with label $2k$ for $k>1$. As all these elements lie in $A_{\operatorname{st}_{\Gamma}(a)}$, so does the subgroup $C_{G}(a)$.
@@
(2) $z_{a,w}=a^{1-k}(wa)^{k}a^{-1}=w^{a^{k-1}}w^{a^{k-2}}\ldots w,$
for each $w\in V$ linked with $a$ with label $2k$ for $k>1$ (here we use the convention that $w^{a}=a^{-1}wa$).
As $A_{\operatorname{st}_{\Gamma}(a)}$ is parabolic, we have an explicit presentation for $A_{\operatorname{st}_{\Gamma}(a)}$. The generators are the vertices of $\operatorname{st}_{\Gamma}(a)$ and there are two types of relators that involve $a$: some of the form $[v,a]$ and some of the form $(wa)^{k}=(aw)^{k}$ for $k>1$.
We can modify this presentation using Tietze transformations as follows. We add generators of the form $w_{i}=w^{a^{i}}$ for $21\leq i\leq k-$ 2 and the elements $z_{a,w}$ as above. Writing $w_{0}$ instead of $w$, we get new relators
(3) $w_{i}^{a}=w_{i+1}$
@@
and furthermore, the relator $(wa)^{k}=(aw)^{k}$ yields
@@
Therefore, we get a new presentation as follows. As generators, we have $a$ together with the set $\Omega_{1}$ of vertices $v$ in $\operatorname{lk}_{\Gamma}(a)$ such that the label of the edge $\{a,v\}$ is 2 and with the set $\Omega_{2}$ consisting of the $k$ elements
$w_{0},w_{1},\ldots,w_{k-2},z_{a,w}$
for each vertex $w$ in $\operatorname{lk}_{\Gamma}(a)$ such that the label of the edge $\{a,w\}$ is $2k$.
The relators that involve $a$ are of the following form. For each $v\in\Omega_{1}$, $v^{a}=v$, and for each $w$ linked to $a$ with label $2k$, the relators in equation (3) and the relator $z_{a,w}^{a}=z_{a,w}$.
To understand which are the relators that do not involve $a$, recall that the FC-condition implies that if two vertices $w$ and $u$ are linked to $a$ with labels greater than 2, then they cannot be adjacent. If such a $w$ is linked to some $v\in V$ with $[v,a]=1$, then the label must be 2, so we also get that $[w_{i},v]=1$ for $20\leq i\leq k-$ 2 and $[z_{a,w},v]=1$. Moreover, we have the relators given by the subgraph of $\Gamma$ with $\Omega_{1}$ as vertex set.
Now, let $M$ be the abstract group defined by a presentation having as generators the elements in $\Omega_{1}$ and $\Omega_{2}$ together with all the previous relators that do not involve $a$. We may define an action of $a$ by conjugation on this group using the relators that do involve $a$. To check that this action is well-defined we argue as follows. The only instances in which it may happen that two generators $x,y$ of $M$ are related (in $M$) are when either both lie in $\Omega_{1}$, and in that case the $a$-action leaves both invariant, or if one, say $x$, lies in $\Omega_{1}$ and the other, say $y$, is one of the $w_{i}$’s or $z_{a,w}$. In this last case the FC-condition mentioned above implies that the relation between $x$ and $y$ must be of type $[x,y]=1$ and the definition of the $a$-action implies that in all the possible cases we have that $[x^{a},y^{a}]=1$.
Therefore, we can define the (abstract) semidirect product $\langle a\rangle\ltimes M$. But the standard semidirect product presentation for this group is precisely the presentation for $A_{\operatorname{st}_{\Gamma}(a)}$ that we had before, so both groups are isomorphic. In fact, looking at the presentation given for $M$, we see that it is the Artin group associated to the graph $\hat{\Gamma}$ with vertices precisely the elements in $\Omega_{1}\cup\Omega_{2}$ and edges determining the relators of $M$. Let $L\leq M$ be the subgroup generated by the subgraph determined by $\Omega_{1}$ and all the vertices of the form $z_{a,w}$ of the $\Omega_{2}$ set. Then, $L$ together with $a$ generate $C_{G}(a)$ (see equation (2) and the paragraph before). Moreover, the previous construction implies that $L$ is the Artin group associated to a subgraph of $\hat{\Gamma}$ which is isomorphic to $\operatorname{lk}_{\Gamma}(a)$ (but with different vertex labels).
@@
We have that $C_{G}(a)\leq A_{\operatorname{st}_{\Gamma}(a)}$ (see Lemma 5.1), so
@@
Consider first the case when $S=\{a\}$ has only one element. Then, as $N_{G}(A_{S})=C_{G}(a)$ (see (1)), using Lemma 5.1 we have (iii).
@@
where the first equality is (1). Using the previous case we get again (iii).
@@
For this, it is enough to show that $C_{G}(A_{S})=C_{G_{1}}(A_{S})$ (recall (1)). Consider the subgraph induced by the three vertices $a$, $b$, $v$. Lemma 5.1 implies that $C_{G}(a)\cap C_{G}(b)=C_{G_{1}}(a)\cap C_{G_{1}}(b)$. Therefore,
@@
This implies that we can successively remove from $\Gamma$ vertices of $\Gamma\setminus S$ without affecting the normalizer until we get a subgraph $S\subseteq\Gamma_{2}\subseteq\Gamma$ such that for any vertex $w$ of $\operatorname{lk}_{\Gamma_{2}}(S)$ and any $a\in V(S)$ so that $w$ and $a$ are linked by an edge of label $2k$ with $k>1$, the vertex $a$ commutes with any other vertex in $S$, that is, it is central in $A_{S}$. Hence, there is some $Z\subseteq S$ (possibly empty) with $A_{Z}$ central in $A_{S}$ and such that all the edges between $S$ and $\operatorname{lk}_{\Gamma_{2}}(S)$ with label greater than 2 have an endpoint in $Z$. As $S$ is not free abelian, $\varnothing\neq S\setminus Z$, and the elements of $A_{S\setminus Z}$ commute with the elements of $A_{\operatorname{lk}_{\Gamma_{2}}(S)}$. Therefore, for $G_{2}=A_{\Gamma_{2}}$ we have that
@@
We follow the steps taken by Minasyan-Osin in [25, Lemma 6.10] where they show the analogous result for graph products of groups.
@@
where we abuse notation and we denote by $S(Y)$ the subgraph of $\Gamma$ induced by $S(Y)$. Since $\Gamma$ is finite, the function, that assigns to each finite subset $Y\subseteq X$ the integer value $|S(Y)|$, attains its maximum on some finite subset $X^{\prime}\subseteq X$. Hence, to prove the result it is enough to show that $X\subseteq P_{c_{\Gamma}}(X^{\prime})$.
@@
$P_{c_{\Gamma}}(X^{\prime})=P_{c_{\Gamma}}(H).$
The subgroup $H^{\prime}\coloneqq\langle X^{\prime}\rangle\leq H$ is finitely generated and every element of $H^{\prime}$ fixes a vertex of $T$, so $H^{\prime}$ fixes some vertex of $T$ (see, for instance, [29, (3.4)]). Hence, there is $f\in G$ such that $H^{\prime}\leq fA_{D}f^{-1}$, where $D$ is either $\operatorname{st}_{\Gamma}(w)$ or $\Gamma\setminus\{w\}$. It follows that
@@
    there are Artin groups $A_{T_{1}}$, $A_{T_{2}}$ with $0<|V(T_{1})|,|V(T_{2})|<|V(\Gamma)|$ such that $H\leq A_{T_{1}}\times A_{T_{2}}$; or
@@
The proof is similar to the one of [25, Corollary 6.19]. Hence, we refer the reader to such paper for a deeper understanding of the main concepts and results.
@@
Let us secondly prove that we can find a hyperbolic element $g\in H$ such that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is normalized by $H$ in $A_{\Gamma}$. Following [25, Definition 6.2], if $P=fA_{S}f^{-1}$ is a parabolic subgroup of $A_{\Gamma}$ for some subgraph $S$ of $\Gamma$ and $f\in A_{\Gamma}$, then the parabolic dimension $\text{pdim}_{\Gamma}(P)$ of $P$ is the number of vertices of $S$. Then [1, Lemma 3.2] ensures that $\text{pdim}_{\Gamma}(P)$ is well-defined, as the result shows that if $g_{1}A_{S}g_{1}^{-1}\subseteq g_{2}A_{T}g_{2}^{-1}$ for some subgraphs $S,T$ of $\Gamma$ and $g_{1},g_{2}\in A_{\Gamma}$, then $S\subseteq T$.
We claim that if $g$ is a hyperbolic element such that $\text{pdim}_{\Gamma}(\text{PStab}_{A_{\Gamma}}(\mathcal{L}))$ is minimal, then $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is normalized by $H$ in $A_{\Gamma}$. This is exactly shown as [25, Proposition 6.13]. The properties used here are proved in [1], and they correspond to [1, Lemma 3.1, Lemma 3.2, Lemma 3.4].
By Theorem 5.3 there are indeed hyperbolic elements in $H$ for the given action on the Bass-Serre tree $T$. If we consider $h$ an element in $H$ which is hyperbolic and such that $\text{pdim}_{\Gamma}(\text{PStab}_{A_{\Gamma}}(\mathcal{L}))$ is minimal, where $\mathcal{L}$ is the axis of $h$, then we have shown that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is normalized by $H$ and parabolic in $A_{\Gamma}$, so there is a subgraph $S$ of $\Gamma$ and $f\in A_{\Gamma}$ such that $fA_{S}f^{-1}$ is normalized by $H$. But then $H\leq N_{G}(fA_{S}f^{-1})$ and as $H$ is not a subgroup of $fA_{S}f^{-1}$ (otherwise $H$ would fix a point in $T$) by Proposition 5.2, we get (i). Otherwise, we conclude that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})$ is trivial. We have shown that there are two vertices $x$ and $y$ such that $\text{PStab}_{A_{\Gamma}}(\mathcal{L})=\text{PStab}_{A_{\Gamma}}(\{x,y\})$, so [25, Corollary 4.3] implies that $h$ satisfies the WPD condition. Then, $A_{\operatorname{lk}_{\Gamma}(w)}$ is weakly malnormal, so from [25, Corollary 2.2], it follows that the action of $H$ on the associated Bass-Serre tree is acylindrically hyperbolic (since we are assuming that $H$ is not virtually cyclic). ∎
@@
If $H$ is a subgroup of $A_{\Delta}$ such that $\Delta$ is a complete graph, then either we can embed $H$ in a decomposition as before or $A_{\Delta}$ is dihedral (recall that $\Delta$ is even of FC-type). If $G=A_{\Delta}$ is dihedral associated to the edge $\{a,b\}$ with label $2k$, the defining relator is $(ab)^{k}=(ba)^{k}$. Setting $c$ to be $ab$, this relator can be transformed into $c^{k}b=bc^{k}$. Then, if we denote by $C$ the subgroup generated by $c^{k}$, we see that $C$ is central in $G$ (in fact, it is the center of $G$). Let $F$ be the normal subgroup of $G$ normally generated by $b$. It is known (see, for example, [2, Lemma 2.5]) that $F$ is free and $C\times F$ has finite index in $G$. Moreover, using the presentation of $G$ in terms of $c$ and $b$ we have that $G/C=\mathbb{Z}\ast\mathbb{Z}_{k}$. Now, if $H\cap C=1$, then $H$ is isomorphic to a subgroup of $G/C$ and by [25, Theorem 2.12] we conclude that $H$ is acylindrically hyperbolic. Thus, we assume that $H\cap C\neq 1$. But as $H$ is not virtually cyclic, $H\cap F\neq 1$ and we are in case (i).
@@
Recall that $N_{1}$ and $N_{2}$ above must be infinite because even Artin groups of FC-type are torsion-free. Acylindrically hyperbolic groups cannot contain infinite normal subgroups as in (1) of Corollary 5.7 (see [25, Theorem 3.7 (b), (c)]), so this result implies that for subgroups of even Artin groups of FC-type this obstruction characterizes acylindrical hyperbolicity.